# Strategic Informed Trading and the Value of Private Information

### v1

@@ 1. The Equilibrium
The no-signal equilibrium corresponds to $p_{I}=0$. In the price-taking case, the equilibrium price and the risk aversion adjusted optimal positions are $p_{ns}(Z_{N})$ and $\widehat{\psi}_{ns,I}=\widehat{\psi}_{ns,U}=\widehat{\psi}_{ns}(Z_{N})$ where (recall (2) and (19))
(28) $p_{ns}(z)=p_{0}+\lambda\frac{z}{\alpha_{I}};\qquad\widehat{\psi}_{ns}(z)=\widehat{\Pi}-\lambda\frac{z}{\alpha_{I}},$
where $\lambda$ is from (17). In the price-impact case, the equilibrium price is $p_{ns,\iota}(Z_{N})$ and the risk aversion adjusted optimal policies for $I$ and $U$ are $\widehat{\psi}_{ns,\iota,I}(Z_{N})$ and $\widehat{\psi}_{ns,\iota,U}(Z_{N})$, where
$\begin{split}p_{ns,\iota}(z)&=p_{0}+\frac{\lambda}{1-\lambda^{2}}\frac{z}{\alpha_{I}};\qquad\widehat{\psi}_{ns,\iota,I}(z)=\widehat{\Pi}-\frac{\lambda}{1+\lambda}\frac{z}{\alpha_{I}};\qquad\widehat{\psi}_{ns,\iota,U}(z)=\widehat{\Pi}-\frac{\lambda}{1-\lambda^{2}}\frac{z}{\alpha_{I}}.\end{split}$
@@ 2. Comparison Analysis: Signals and Price Sensitivity
In this section, we compare the public signals and price sensitivity with respect to signals of the two equilibria. We label the price impact equilibria as “PI” and the price taking equilibria as “PT”. We show the PI public signal is of a worse quality, and prices are less responsive to not only the market and insider signals, but also to the publicly observable (risk-tolerance weighted) insider’s and noise trader residual demand $\widehat{\psi}_{I}-\psi_{I,0}+Z_{N}/\alpha_{I}$. Thus, the main message of this section is
@@
Throughout, we include the subscript $\iota$ when describing any quantity obtained internalizing price impact. Proofs of all results are in Appendix C. Lastly, using (17), we see that $p_{U,\iota}$ from (13) and $p_{U}$ from (23) take the form
@@
As we have seen, in both the PI and PT equilibria a signal of the form “$X+\textrm{Noise}$” is communicated to market. It is natural to ask which signal is of a higher quality, or even more pointedly, is the public signal less informative in the presence of price impact? To address these questions we write the market signals as functions of the insider signal $G$ and noisy demand $Z_{N}$, and using (11), (12) and (22) we obtain
@@
From Theorem 1.6 we know $\widehat{y}>0$, which implies
@@
That the public signal is less informative under price impact is associated with the way the uniformed trader determines his demand. Indeed, because the uniformed trader accounts for the insider’s internalization of price impact, in his optimization problem he considers $p_{U,\iota}$ instead of $p_{U}$ (see the demand functions (15) and (25)). In other words, he recognizes the insider reveals a wangled signal, and responds with a less elastic demand function.
Remark 2.2.
Accounting for price impact changes the public signal in the direction of the noise traders’ (liquidity providers’) order. As $\widehat{y}>0$, we have $(h_{\iota}(g,z)-h(g,z))z\geq 0$ for all $(g,z)$. In words, the strategically revealed signal by the insider is higher if and only if there is positive demand from the noise traders. As we will analyze in Section 4, positive $z$ implies the insider and uniformed trader take short positions at equilibrium, and hence the strategically enhanced public signal increases the price that the traders sell the assets to the liquidity providers.
@@
In both the PI and PT equilibria, prices are affine functions of the respective public signals. However, the coefficients in the functions differ. This leads one to ask whether price impact increases or decreases the sensitivity of prices with respect to public signaling. To answer this, we first re-express the prices from (20) and (24) using the notation of (17) and (29).
@@
(31) $\begin{split}p_{\iota}(h_{\iota})&=p_{0}+\frac{p_{I}\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\left(h_{\iota}-p_{0}\right);\\ p(h)&=p_{0}+\frac{p_{I}(\kappa p_{I}+\lambda)}{1-\lambda+(1+p_{I})(\kappa p_{I}+\lambda)}\left(h-p_{0}\right).\end{split}$
Given this, we now consider reactivity to the market and, equivalently in view of (30), insider signals. To do so, define the slopes
@@
The following proposition shows that prices are always more reactive to the insider, and hence to the market signal, when the insider does not internalize price impact. This is directly linked with the lower elasticity of the uniformed trader’s demand function due to price impact. Note that this is an endogenously derived outcome. The insider has a motive to make the public signal noisier, which in turn makes uniformed trader less elastic and yields prices which are less sensitive to the public signal.
@@
Using (9), (14), (17), (26) and (29), prices are affine in the combined demand with respective slopes
@@
The equilibrium prices is less sensitive to the publicly observable combined demand in the PI equilibria: $m_{\widehat{\chi},\iota}<m_{\widehat{\chi}}$.
3. Welfare Analysis
Overview
This section is dedicated to analyzing the traders’ welfare. We primarily study two issues: how the insider’s signal quality is translated to her welfare from trading; and the effect of price-impact internalization on traders’ welfare. We again label the price impact equilibrium PI and the price taking equilibrium PT.
Following Laffont (1985), we define welfare at both the ex-ante level (i.e. at time $0-$, prior to signal revelation) and interim level (at time 00, after the signal revelation) in terms of certainty equivalents. As such one can alternatively think of this comparison as a comparison of indirect utility gains from trading. We use the subscripts $0-$ and 00 to indicate ex-ante and interim welfare respectively. Welfare will always be computed using the overall wealth in (5). Given this, we denote by $\widehat{\mathcal{W}}_{I,\iota},\widehat{\mathcal{W}}_{U,\iota}$ the optimal terminal wealths in the PI equilibrium, and $\widehat{\mathcal{W}}_{I},\widehat{\mathcal{W}}_{U}$ those in the PT case. Then, for $k\in\left\{\ ,\iota\right\}$ the corresponding interim certainty equivalents are
$\begin{split}\textrm{CE}^{I}_{0,k}&=-\alpha_{I}\log\left(\mathbb{E}\left[e^{-(1/\alpha_{I})\widehat{\mathcal{W}}_{I,k}}\big{|}\sigma(G,H_{k})\right]\right),\\ \textrm{CE}^{U}_{0,k}&=-\alpha_{U}\log\left(\mathbb{E}\left[e^{-(1/\alpha_{U})\widehat{\mathcal{W}}_{U,k}}\big{|}\sigma(H_{k})\right]\right),\end{split}$
@@ 3. Welfare Analysis
$\textrm{CE}^{j}_{0-,k}=-\alpha_{j}\log\left(\mathbb{E}\left[e^{-(1/\alpha_{j})\widehat{\mathcal{W}}_{j,k}}\right]\right);\qquad j\in\left\{I,U\right\}.$
We now summarize our main findings. First, in the PI equilibrium, insider welfare is monotonically increasing in the precision $p_{I}$. As mentioned in the introductory section, this is not a standard result, as in the majority of private-information models, better precision does not always imply higher utility gains. For instance, this result stands in direct contrast to Nezafat and Schroder (2023), and shows the conclusions there-in are a consequence of assuming the uninformed trader also internalizes his impact on prices. Additionally, this result is also in contrast to Grossman and Stiglitz (1980), where all traders are price takers. Therefore, one only obtains the reasonable conclusion that a better signal is better for the insider (if not, why would the insider expend effort to obtain the signal?) when there is a differential between the insider and uniformed traders, not only in terms of information, but also in terms of their internalization of price impact.
Second, insider ex-ante welfare is not always higher in the PI equilibrium. While typically this is true, as the insider becomes increasingly more risk averse, PT welfare will exceed PI welfare when also the uninformed trader has sufficiently high risk tolerance, and provided the insider signal quality is not too low. However, in contrast to the insider, uninformed ex-ante welfare is always higher in the PI case.
Third, absent private information, insider welfare is always higher in the PI case, and remarkably this holds at the interim level. In fact, when insider and uniformed traders have the same risk tolerance we can order interim welfare as follows
@@
so that uninformed trader’s welfare exceeds insider’s welfare in the price-impact case. We use the notation in (17) and, aside from Proposition 3.4, all proofs in this section are in Appendix E.
@@
We start by calculating the certainty equivalents. Propositions D.4 and D.5 explicitly compute ex-ante welfare for both types of traders in the PI and PT equilibria under the general model of Appendix A. While the formulas there-in are very long in the generalized model, a significant simplification occurs in the model of Section 1 and using the notation in (17). To state the proposition define
@@
as the certainty equivalents absent private information for the initial allocations in (2).
@@
Welfare and the insider’s signal precision
Here we analyze insider welfare with respect to the signal precision $p_{I}$. Though not modeled, it presumably costs effort, time and/or money to both obtain and refine the signal. In fact, one of the most important questions in the related literature is whether it is worth for the traders who have the ability and resources to obtain a private signal to actually pay the cost and obtain it. As such, one should examine whether the benefits of the private signal (as measured by welfare) are increasing with respect to the quality of a signal (as measured by signal precision). This is of course connected with the related cost, in the sense that better signal is normally linked to higher cost.
@@
where $\widehat{y}=\widehat{y}(p_{I})$ is the unique positive solution of (18). Numerically, this can easily seen to be increasing in $p_{I}$ by randomly sampling $\kappa>0$, $\lambda\in(0,1)$, solving for $\widehat{y}(p_{I})$ and then plotting $p_{I}\to\phi_{\iota}(p_{I})$. However, we offer an analytic proof in the following proposition.
Proposition 3.2.
@@
On the other hand, for the price-taking equilibrium, we have the map
@@
When $\lambda\leq 1/2$ (equivalently $\alpha_{I}\leq\alpha_{U}$ or that insider is less risk tolerant than uniformed traders), $\phi$ is increasing at 00. However, when the insider is relatively more risk tolerant ($\lambda>1/2$), $\phi$ will be decreasing at 00 for $\kappa=\alpha_{I}^{2}p_{N}$ small enough (which can happen if the noise trader variance/volume is very large).
Interestingly enough, there are cases where the certainty equivalents with and without price impact have the opposite monotonicity with respect to signal’s precision. This is pictured in Figure 1 which shows when insider does not internalize her price impact the better quality of her signal, the lower her ex-ante expected utility becomes; while under price impact we have the more reasonable situation where price impact materializes the better quality of the signal to higher insider’s certainty equivalent.
Remark 3.3.
The above discussion highlights a very interesting feature of the price-impact model. If the insider does not internalize her price impact, her welfare created by the private signal is not necessarily increasing in the signal’s quality. This means that in the PT equilibrium, it is not always worth it to obtain a better signal, as it may decrease utility. On the other hand, when she does internalize her price impact (and the other traders are price-takers), it is always beneficial for the insider to try and improve the quality of her private signal, as long as this improvement comes with marginally lower cost than the corresponding increase of certainty equivalent [^12] Note that beneficial price-impact for the insider does not mean that the uniformed traders suffer loss of utility because of price impact. On the contrary, when uniformed traders are the same side of trade with the insider, not only they gain welfare due to price impact, but their gain may be even higher than the insider’s (e.g. under absence of private signal, see Proposition E.1 and Remark 3.8 below)..
@@
Price-impact and price-taking welfare comparison
Next, we focus on how price impact internalization affects both informed and uniformed trader’s welfare. In particular, we examine whether the internalization of price impact implies higher welfare for the insider and the uniformed trader, when compared to the price taking case. Using Proposition 3.1 we readily get the relations
(35) $\begin{split}\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}&\quad\Longleftrightarrow\quad\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{(1+p_{I})(1+2\widehat{y})}\geq\frac{(1-\lambda)^{2}\kappa p+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}};\\ \textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}&\quad\Longleftrightarrow\quad\frac{\kappa p_{I}+(1+\widehat{y})^{2}}{(1-\lambda)^{2}(1+2\widehat{y})^{2}}\geq\frac{1+\kappa p_{I}}{(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}.\end{split}$
Our first result shows that insider welfare need not increase when she internalizes her price impact.
@@
Numerically, this is demonstrated in Figure 2. Indeed, in the joint combination of high $\alpha_{U}$ (uninformed close to risk neutrality) and low-to-moderate $p_{I}\in(0,2)$ (modest signal quality) welfare may decrease. Analytically, this will be shown in Proposition 3.6. ∎
Refer to captionRefer to captionRefer to caption
@@
As already mentioned, we give the economic intuition behind these model’s predictions in Section 4. For now, we state a couple of additional results which will clarify more the situation.
@@
It is rather complicated to precisely describe the set of input parameters that characterizes the order of insider’s certainty equivalents $\left\{(\kappa,p_{I},\lambda)\ |\ \textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}\right\}$. From Figure 2, we conjugate however that the situation becomes clearer when we consider the asymptotics with respect to traders’ risk tolerances. For this, in the following proposition, we take limits as the insider and uninformed traders’ risk tolerance go to 00 and $\infty$.
@@
Fix $p_{I},p_{N}$. Then,
  1. (1)
    Fix $\alpha_{I}$. As $\alpha_{U}\to 0$, $\textrm{CE}^{I}_{0-}$ remains bounded while $\textrm{CE}^{I}_{0-,\iota}\to\infty$. As $\alpha_{U}\to\infty$,
    $\lim_{\alpha_{U}\to\infty}\left(\textrm{CE}^{I}_{0-,\iota}-\textrm{CE}^{I}_{0-}\right)>0.$
  2. (2)
    Fix $\alpha_{U}$. As $\alpha_{I}\to 0$, $\lim_{\alpha_{I}\to 0}(\textrm{CE}^{I}_{0-,\iota}-\textrm{CE}^{I}_{0-})/\alpha_{I}=0$ but
@@
    Thus, for $\alpha_{I}$ small, $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ if and only if $1+p_{I}-\alpha_{U}^{2}p_{I}p_{N}\geq 0$. As $\alpha_{I}\to\infty$,
    $\lim_{\alpha_{I}\to\infty}\left(\textrm{CE}^{I}_{0-,\iota}-\textrm{CE}^{I}_{0-}\right)>0.$
  3. (3)
    If $\alpha_{I}=\alpha_{U}$, then $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$.
Proposition 3.6 implies the risk aversion is a crucial parameter for the certainty equivalents’ comparisons. Indeed, when traders have the same risk aversion internalizing price impact always increases agents’ welfare. On the other hand, fixing $\alpha_{U},p_{N}$, insider welfare may be lower than in the price-taking case when she is very risk averse and the following structure condition holds
@@
Welfare in the absence of private information
We conclude our welfare analysis proving that if one turns off the asymmetric information channel, and focuses solely on effects due to internalizing of price impact, then remarkably, insider welfare increases at the interim level. As mentioned above, the situation with no private signal assumes only the insider possesses and exploits her price impact. The other (non-noise) traders act as price takers. This corresponds to when a (large) risk averse agent acts strategically, even when she does not have access to a private signal, and when the uninformed traders represent the mass of all other rational agents who are relatively small (i.e. possess no market power) and who also have no private signal. In sum, we are measuring the effects of assuming heterogeneity, not in information, but in internalization of price impact.
For this, recall from Proposition 1.11 that the no-signal equilibrium corresponds to taking $p_{I}\to 0$. Using the interim welfare formulas in Lemmas D.1 and D.2 in Appendix D, we obtain the associated no-signal equilibrium quantities in the following proposition. To state it, recall $\lambda$ from (17).
@@
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{I}_{0}(G,Z_{N})&=\textrm{CE}_{nsn}^{I}+\frac{\lambda^{2}}{2\alpha_{I}}Z_{N}^{2},\qquad\lim_{p_{I}\to 0}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})=\textrm{CE}_{nsn}^{I}+\frac{\lambda^{2}}{2(1-\lambda^{2})\alpha_{I}}Z_{N}^{2}.\end{split}$
Therefore, insider interim welfare always increases when internalizing price impact. For the uninformed agent we obtain almost surely
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{U}_{0}(H)&=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{U}\lambda^{2}}{2\alpha_{I}^{2}}Z_{N}^{2},\qquad\lim_{p_{I}\to 0}\textrm{CE}^{U}_{0,\iota}(H_{\iota})=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{U}\lambda^{2}}{2\alpha_{I}^{2}(1-\lambda^{2})^{2}}Z_{N}^{2},\end{split}$
and hence interim welfare also always increases when the insider internalizes price impact.

### v2

@@ 1. The Equilibrium
No-signal equilibria correspond to $p_{I}=0$. In the price-taking case, the equilibrium price and optimal positions are $p_{ns}(Z_{N})$ and $\widehat{\psi}_{ns,I}(Z_{N})=\widehat{\psi}_{ns,U}(Z_{N})=\widehat{\psi}_{ns}(Z_{N})$ where (recall (1), (7) and (8))
(20) $p_{ns}(z)=p_{{\mathbb{Q}}_{0}}+\lambda\frac{z}{\alpha_{I}};\qquad\widehat{\psi}_{ns}(z)=\widehat{\Pi}-\lambda\frac{z}{\alpha_{I}}.$
In the price-impact case, the equilibrium price is $p_{ns,\iota}(Z_{N})$ and the optimal policies are $\widehat{\psi}_{ns,I,\iota}(Z_{N})$ and $\widehat{\psi}_{ns,U,\iota}(Z_{N})$, where
$\begin{split}p_{ns,\iota}(z)&=p_{{\mathbb{Q}}_{0}}+\frac{\lambda}{1-\lambda^{2}}\frac{z}{\alpha_{I}};\quad\widehat{\psi}_{ns,I,\iota}(z)=\widehat{\Pi}-\frac{\lambda}{1+\lambda}\frac{z}{\alpha_{I}};\quad\widehat{\psi}_{ns,U,\iota}(z)=\widehat{\Pi}-\frac{\lambda}{1-\lambda^{2}}\frac{z}{\alpha_{I}}.\end{split}$
A comment on the linearity of price impact
Our assumption of linear price impact is motivated by the price-taking case. Indeed, using (17), (18) one can show
(21) $\begin{split}p(h(g,z))&=p_{{\mathbb{Q}}_{0}}+\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{U}(p_{I}-p_{U})}\left(\widehat{\psi}_{I}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}\right).\end{split}$
This is the reverse combined demand function at equilibrium, and indicates linear price impact. Indeed, even though the insider does not internalize impact in the price-taking case, in equilibrium the price is linearly impacted by her trade, combined with the noise trader’s demand. The price takes the form (5), where
$V_{p}=p_{{\mathbb{Q}}_{0}}=-\widehat{\Pi},\qquad M_{p}=\frac{\alpha_{I}p_{I}+\alpha_{U}p_{U}}{\alpha_{U}(p_{I}-p_{U})}.$
As discussed in the introduction, as long as the insider internalizes her price impact and the uninformed trader takes this into account, the price-taking and price-impact equilibria cannot coincide. In fact, even if the insider submits the demand which is optimal in the price-taking equilibrium, if she internalizes price impact, the market will not equilibrate to the price-taking equilibrium price. This would be the case if the uninformed trader did not perceive the change in market signal precision due to the insider’s demand (i.e., if he ignored the insider’s internalization of the price impact and assumed $p_{U,\iota}(y)=p_{U}$ for all $y$.)
On the other hand, there is a $y^{*}>0$ such that $M_{p,\iota}(y^{*})=y^{*}/(1+p_{I})$ coincides with $M_{p}$ from above. For this $M_{p,\iota}(y^{*})$, if the insider used the price-taking optimal demand $\widehat{\psi}_{I}$ from (18), it would reveal the same signal to the uninformed trader as in the price-taking equilibrium. This would lead to the same uninformed trader’s demand and hence the same clearing price as in price-taking equilibrium. However, when $M_{p,\iota}=M(y^{*})$ and $V_{p,\iota}=-\widehat{\Pi}$, $\widehat{\psi}_{I}$ from (18) is not optimal for the insider in the price-impact model, hence the price-taking and price-impact equilibria cannot be the same.
@@ 2. Comparison Analysis: Signals and Price Sensitivity
In this section, we compare the public signals and price sensitivity with respect to signals of the two equilibria. As in the introduction, we label the price-impact equilibria as “PI” and the price-taking equilibria as “PT”. We show the PI public signal is of a worse quality, and prices are less responsive to not only the market and insider signals, but also to the publicly observable (risk-tolerance weighted) insider’s and noise trader residual demand $\widehat{\psi}_{I}-\psi_{I,0}+Z_{N}/\alpha_{I}$. Thus, the main message of this section is
@@
Throughout, we include the subscript $\iota$ when describing any quantity obtained internalizing price impact. Proofs are in Appendix B.3, and we collect $p_{U,\iota}$ from (10) and $p_{U}$ from (16)
@@
As we have seen, in both the PI and PT equilibria a signal of the form “$X+\textrm{Noise}$” is communicated to market. It is natural to ask which signal is of a higher quality, or even more pointedly, is the public signal less informative in the presence of price impact? To address these questions we write the market signals as functions of the insider signal $G$ and noisy demand $Z_{N}$, and using (9), (15) we obtain
@@
From Theorem 1.2 we know $\widehat{y}>0$, which from (22) implies
@@
That the public signal is less informative under price impact is associated with the way the uninformed trader determines his demand. Indeed, because the uninformed trader accounts for the insider’s internalization of price impact, in his optimization problem he considers $p_{U,\iota}$ instead of $p_{U}$ (compare the demand functions (13) and (19)). In other words, he recognizes the insider reveals a wangled signal, and responds with a less elastic demand function.
@@
In both the PI and PT equilibria, prices are affine functions of the respective public signals. However, the coefficients in the functions differ. This leads one to ask whether price impact increases or decreases the sensitivity of prices with respect to public signaling. To answer this, we recall the PI price $p_{\iota}$ from (11) and re-express the PT price (17) using the notation of (7).
@@
(24) $\begin{split}p_{\iota}(h_{\iota})&=p_{{\mathbb{Q}}_{0}}+\frac{p_{I}\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\left(h_{\iota}-p_{{\mathbb{Q}}_{0}}\right);\\ p(h)&=p_{{\mathbb{Q}}_{0}}+\frac{p_{I}(\kappa p_{I}+\lambda)}{1-\lambda+(1+p_{I})(\kappa p_{I}+\lambda)}\left(h-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
Given this, and in view of (9), (15), define the slopes
@@
The following proposition shows prices are always more reactive to the insider, and hence to the market signal, when the insider does not internalize price impact. This is directly linked with the lower elasticity of the uninformed trader’s demand function due to price impact, and is an endogenously derived outcome. The insider is motivated to make the public signal noisier, which makes uninformed trader less elastic and yields prices which are less sensitive to the public signal.
@@
Using Theorem 1.2, Proposition 1.3 and (7) one can show prices are affine in the combined demand with respective slopes
@@
The equilibrium prices are less sensitive to the publicly observable combined demand in the PI equilibria: $m_{\widehat{\chi},\iota}<m_{\widehat{\chi}}$.
3. Indirect Utility
This section is dedicated to analyzing the traders’ indirect utility (called “utility” or “value”). We address how the insider’s signal quality relates to her utility, and how price-impact internalization effects all agents’ utility. We again label the price impact equilibrium PI and the price taking equilibrium PT.
We define utility at both the ex-ante level (i.e. at time $0-$, prior to signal revelation) and interim level (at time 00, after the signal revelation) in terms of certainty equivalents. Utility will always be computed using the overall wealth in (4). Given this, we denote by $\widehat{\mathcal{W}}_{I,\iota},\widehat{\mathcal{W}}_{U,\iota}$ the optimal terminal wealths in the PI equilibrium, and $\widehat{\mathcal{W}}_{I},\widehat{\mathcal{W}}_{U}$ those in the PT case. Then, for $k\in\left\{\ ,\iota\right\}$ the corresponding interim certainty equivalents are
$\begin{split}\textrm{CE}^{I}_{0,k}&=-\alpha_{I}\log\left(\mathbb{E}\left[e^{-\frac{1}{\alpha_{I}}\widehat{\mathcal{W}}_{I,k}}\big{|}\sigma(G,H_{k})\right]\right);\quad\textrm{CE}^{U}_{0,k}=-\alpha_{U}\log\left(\mathbb{E}\left[e^{-\frac{1}{\alpha_{U}}\widehat{\mathcal{W}}_{U,k}}\big{|}\sigma(H_{k})\right]\right),\end{split}$
@@ 3. Indirect Utility
$\textrm{CE}^{j}_{0-,k}=-\alpha_{j}\log\left(\mathbb{E}\left[e^{-\frac{1}{\alpha_{j}}\widehat{\mathcal{W}}_{j,k}}\right]\right);\qquad j\in\left\{I,U\right\}.$
We find that (i) in the PI equilibrium, insider ex-ante utility is always strictly increasing in the precision $p_{I}$; (ii) insider ex-ante utility is not always higher in the PI equilibrium; and (iii) absent private information, insider utility is always higher in the PI equilibrium, and remarkably this holds at the interim level. In fact, when the insider and uninformed traders have the same risk tolerance we can order interim utility
@@
so the uninformed trader’s utility exceeds the insider’s utility in the PI case. We use the notation in (7) and, Proposition 3.3 aside, all proofs of this section are in Appendix B.5.
@@
We start calculating the certainty equivalents. Propositions B.1 and B.2 (see also Remark B.3) below compute ex-ante utilities in the PI and PT equilibria under the general model of Appendix A. To state the results in the model of Section 1 define
@@
as the certainty equivalents absent private information for the allocations in (1).
@@
Utility and the insider’s signal precision
Though not modeled, it presumably costs effort, time and/or money for the insider to obtain and refine her signal. In fact, one of the central questions in the literature (see among others Kacperczyk and Pagnotta (2019) and Kacperczyk et al. (2024)) is whether traders who have the ability/resources to obtain a private signal should actually pay the cost and obtain it. To answer this question, one should examine whether the benefits of the private signal (as measured by utility) are increasing with respect to the quality of a signal (as measured by signal precision). This is of course connected with the cost, in that a better signal is normally linked to a higher cost.
@@
where $\widehat{y}=\widehat{y}(p_{I})$ is the unique positive solution of (6). Numerically, this is seen to be increasing in $p_{I}$ by randomly sampling $\kappa>0$, $\lambda\in(0,1)$, solving for $\widehat{y}(p_{I})$, and then plotting $p_{I}\to\phi_{\iota}(p_{I})$. However, we offer an analytic proof in the following theorem.
Theorem 3.2.
@@
Alternatively, in the PT equilibrium, one must study the map
@@
When $\lambda\leq 1/2$ (equivalently $\alpha_{I}\leq\alpha_{U}$ so the insider is less risk tolerant than the uninformed traders), $\phi$ is increasing at 00. However, when the insider is relatively more risk tolerant ($\lambda>1/2$), $\phi$ will be decreasing at 00 for $\kappa=\alpha_{I}^{2}p_{N}$ small enough (which can happen if the noise trader variance/volume is very large).
Interestingly, it is possible for the certainty equivalents with and without price impact to have the opposite monotonicity with respect to the signal precision. This is pictured in Figure 2. We conclude that if the insider does not internalize her price impact, her utility may not increase in the signal’s quality. But when she internalizes her price impact, it is always beneficial for her to improve the quality of her private signal, as long as the cost of such improvement does not outweigh the corresponding utility increase.
@@
PI and PT utility comparison
Next, we focus on how price impact internalization affects both informed and uninformed trader’s utility. In particular, we examine whether the internalization of price impact implies higher utility for the insider and the uninformed trader, when compared to the price taking case. Using Proposition 3.1 we readily get the relations
(28) $\begin{split}\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}&\quad\Longleftrightarrow\quad\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{(1+p_{I})(1+2\widehat{y})}\geq\frac{(1-\lambda)^{2}\kappa p_{I}+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}};\\ \textrm{CE}^{U}_{0-,\iota}\geq\textrm{CE}^{U}_{0-}&\quad\Longleftrightarrow\quad\frac{\kappa p_{I}+(1+\widehat{y})^{2}}{(1-\lambda)^{2}(1+2\widehat{y})^{2}}\geq\frac{1+\kappa p_{I}}{(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}.\end{split}$
Our first result shows that insider utility need not increase when she internalizes her price impact.
@@
Numerically, this is demonstrated in Figure 3. Indeed, in the joint combination of high $\alpha_{U}$ (uninformed close to risk neutrality) and low-to-moderate $p_{I}\in(0,2)$ (modest signal quality) utility may decrease. Analytically, this will be shown in Proposition 3.5. ∎
@@
Economic intuition for these results is given in Section 4. Below, we provide explicit comparisons when investor risk aversions become very large or small.
@@
It is rather complicated to precisely describe the parameter set that characterizes the order of insider’s certainty equivalents $\left\{(\kappa,p_{I},\lambda)\ |\ \textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}\right\}$. However, Figure 3 suggests the situation may clarify if we consider asymptotics with respect to traders’ risk tolerances, especially as $\alpha_{I}\to 0$. In the following proposition, we first consider when $\alpha_{I}=\alpha_{U}$ and then take the limit as the insider’s risk tolerance goes to 00.
@@
Fix $p_{I},p_{N}$. If $\alpha_{I}=\alpha_{U}$, then $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$. However, for $\alpha_{U}$ fixed we have
@@
Thus, for small $\alpha_{I}$, $\textrm{CE}^{I}_{0-,\iota}\geq\textrm{CE}^{I}_{0-}$ if and only if $01+p_{I}-\alpha_{U}^{2}p_{I}p_{N}>$ 0.
Proposition 3.5 implies the risk aversion is a crucial parameter for the certainty equivalents’ comparisons. Indeed, when traders have the same risk aversion internalizing price impact always increases traders’ utility. On the other hand, fixing $\alpha_{U},p_{N}$, insider utility may be lower than in the price-taking case when she is very risk averse and the following structure condition holds
@@
Utility absent private information
We conclude our utility analysis proving that if one turns off the asymmetric information channel, and focuses solely on effects due to internalizing of price impact, then insider utility increases at the interim level. As previously mentioned, this setting corresponds to when a (large) risk averse agent acts strategically, even when she does not have access to a private signal, and when the uninformed traders represent the mass of all other rational agents who do not possess market power. In sum, we are measuring the effects heterogeneity in the internalization of price impact.
Recall from Proposition 1.4 that the no-signal equilibrium corresponds to taking $p_{I}\to 0$. Using Proposition B.4 in Appendix B.5, we obtain the following proposition.
@@
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{I}_{0}(G,Z_{N})&=\textrm{CE}_{nsn}^{I}+\frac{\lambda^{2}}{2\alpha_{I}}Z_{N}^{2},\quad\lim_{p_{I}\to 0}\textrm{CE}^{I}_{0,\iota}(G,Z_{N})=\textrm{CE}_{nsn}^{I}+\frac{\lambda^{2}}{2(1-\lambda^{2})\alpha_{I}}Z_{N}^{2}.\end{split}$
Therefore, insider interim utility always increases when internalizing price impact. For the uninformed trader we obtain almost surely
$\begin{split}\lim_{p_{I}\to 0}\textrm{CE}^{U}_{0}(H)&=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{U}\lambda^{2}}{2\alpha_{I}^{2}}Z_{N}^{2},\quad\lim_{p_{I}\to 0}\textrm{CE}^{U}_{0,\iota}(H_{\iota})=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{U}\lambda^{2}}{2\alpha_{I}^{2}(1-\lambda^{2})^{2}}Z_{N}^{2},\end{split}$
and hence interim utility also always increases when the insider internalizes price impact.